Lilian9667:

What is the value of a? a–2=3+6a3 Enter your answer in the box. a =

angiesyl:

a-2=3 6a/3 a-2= 3.3+6a/3 a-2=9+6a/3 3a-6=9+6a 6a-3a= -6-9 3a= -15 a= -5

angiesyl:

hope that helped

Lilian9667:

Thanks SO much